OpenStudy (anonymous):

X2 - x – 2 = 0 a=1 b= -1 c= -2 But then: d = b2 – 4ac = ???

OpenStudy (lalaly):

\[(-1)^2 -4(1)(-2) = 1-(-8) = 9\]